OpenStudy (johnweldon1993):

Motion along a curve, projectiles. The wall in Fenway Park is 37 feet high and 315 feet away from home plate. A baseball hit 3 feet above the ground at an angle 22.5 will just go over is V_o = ______. The time to reach the wall is t = _____ . Check my solution please?

OpenStudy (johnweldon1993):

\[\large x(t) = (v_o cos \alpha)t\] \[\large y(t) = (v_o sin \alpha)t - \frac{1}{2}gt^2\] We want x(t) to be 315 so \[\large 315 = (v_o cos(22.5))t\] solve that for v_o \[\large v_o = \frac{315}{tcos(22.5)}\] sub that in for \(\large v_o\) in y(t) knowing we want y(t) to be ????37??? or ???34??? I did 37 for this \[\large 37 = ((\frac{315}{tcos(22.5)}) sin (22.5))t - \frac{1}{2}(9.8)t^2\] Simplify a bit \[\large 37 = 315tan(22.5) - \frac{1}{2}(9.8)t^2\] \[\large 37 - 315tan(22.5) = - \frac{1}{2}(9.8)t^2\] \[\large 2(37 - 315tan(22.5)) = -9.8t^2\] \[\large t = \sqrt{\frac{2(37 - 315tan(22.5))}{-9.8}}\] \[\large t = 4.37\] And I solve that for velocity etc and I get an answer that checks out for the height being 37 HOWEVER, am I supposed to use the fact that the height started out at 3 feet off the ground and only had to travel 34 feet?

geerky42 (geerky42):

You can use \(x(t)\) to find how long it will take ball to reach the wall, which is 315 feet away from home plate.